OpenStudy (brandonpayan):

Solve for x. 3(3x - 1) + 2(3 - x) = 0

OpenStudy (lovelarap):

first, open the brackets: 9x-3+6-2x=0

OpenStudy (lovelarap):

now put your x together and your simple numbers together

OpenStudy (brandonpayan):

x = 3/7 x = -3/7 x = 1/3 x = -1/3

OpenStudy (brandonpayan):

7x-3?

OpenStudy (lovelarap):

almost, whats -3+6?

OpenStudy (brandonpayan):

3

OpenStudy (lovelarap):

so you have: 7x+3=0 now solve for x

OpenStudy (brandonpayan):

x=7/-3?

OpenStudy (lovelarap):

almost: 7x=-3

OpenStudy (lovelarap):

divide both sides by 7

OpenStudy (bhumi273):

3(3x-1)+2(3x-1)=0 9x-3+6x-2=0 15x-5=0 15x=5 x=5/15 x=1/3